FT-IR MEASUREMENTS OF CROSS SECTIONS OF COLD C_{3}H_{8} IN THE 7 - 15 μm FOR TITAN

\underline{KEEYOON SUNG}, GEOFFREY C. TOON, LINDA R. BROWN, Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Dr.,Pasadena, CA 91109; ARLAN W. MANTZ, Dept. of Physics, Connecticut College, New London, CT 06320; MARY ANN H. SMITH, Science Directorate, NASA Langley Research Center, Hampton, VA 23681.

To support atmospheric remote sensing of Titan, the absorption cross sections of N_{2}-broadened C_{3}H_{8} were obtained at temperatures between 145 and 296 K. For this, 35 spectra of pure- and N_{2}-broadened propane were recorded in the 670 to 1900 cm-1 region using a Fourier transform spectrometer (Bruker IFS-125HR) at the Jet Propulsion Laboratory. A 20.38 cm path temperature-stabilized cryogenic absorption cell was used, which was developed at Connecticut College and described previously [1]. We report the absorption cross sections at the various cold temperatures for nine strong fundamental bands (ν_{26}, ν_{8}, ν_{21}, ν_{20}, ν_{7}, ν_{19}, ν_{18}, ν_{4}, ν_{24}) as well as for many contributions from hot and combination bands. In addition, we present results from ’pseudo -line generation’ (http://mark4sun.jpl.nasa.gov/data/spec/Pseudo/Readme), which includes mean intensities and effective lower state energies on a 0.005 cm^{−1} frequency grid determined in the 690 - 1536 cm^{−1} region from all 35 high-resolution laboratory spectra. It was observed that the pseudo lines reproduce all the observed spectral transmittances well within 3% and the C_{3}H_{8} amounts within 4% on the average. The measured cross sections and synthetic spectra from the pseudoline compilation are compared to earlier work, including the C_{3}H_{8}+N_{2} spectra recorded at PNNL [2] and line-by-line predictions available [3, 4].^{a} ^{b}

a[1] K. Sung, A. W. Mantz, M. A. H. Smith, et al., J Mol Spectrosc 262, 122, 2010.; [2] S. W. Sharpe, et al., Appl Spectrosc 58, 1452, 2004.; [3] J. M. Flaud et al., Mol Phys 108, 699, 2010.; [4] J. M. Flaud et al., J Chem Phys 114, 9361, 2001.

bThe research described in this paper was performed at the Jet Propulsion Laboratory, California Institute of Technology and at The College of William and Mary under contracts with National Aeronautics and Space Administration. US Government Support Acknowledged.